### 11 Apr 2026 v2

# ML Defender (aRGus NDR): An Open-Source Embedded ML NIDS for Botnet and Anomalous Traffic Detection in Resource-Constrained Organizations

Abstract
Ransomware and DDoS attacks disproportionately impact hospitals, schools, and small organizations that cannot afford enterprise security solutions. Existing open-source alternatives either rely on signature matching — missing novel variants — or require infrastructure beyond the reach of their target users. We present ML Defender (aRGus NDR), an open-source network intrusion detection system built in C++20 that brings embedded machine learning inference to organizations with limited resources, deployable on commodity bare-metal hardware at approximately 150–200 USD fully dedicated.
ML Defender implements a six-component pipeline over eBPF/XDP packet capture, ZeroMQ transport, and Protocol Buffers serialization, terminating in a dual-score detection architecture combining a rule-based Fast Detector with an embedded Random Forest classifier. The Maximum Threat Wins policy selects the arithmetic maximum of both detector scores, using ML inference to suppress false positives generated by the heuristic layer.
[…]
We further characterize the pipeline’s throughput ceiling under progressive load in a virtualized environment. Requesting up to 100 Mbps via tcpreplay against the VirtualBox NIC, the pipeline processes up to ${\sim}34$–38 Mbps (the emulated NIC ceiling) with zero packet drops and zero errors across 2.37 million packets. At peak load, ml-detector consumes approximately 3.2 CPU cores; total pipeline RAM remains stable at ${\sim}1.28$ GB with negligible drift (${\sim}18$ MB over 8 minutes). The bottleneck is VirtualBox NIC emulation, not pipeline logic. Post-replay drain behavior confirms that ZeroMQ queues absorb burst traffic correctly and drain asynchronously, validating the queue stability design. All throughput figures should be treated as conservative lower bounds; bare-metal characterization remains future work (§11).
These results demonstrate architectural feasibility under controlled replay conditions rather than universal detection capability. The evaluation covers a single botnet scenario from 2011 with synthetic-trained classifiers; generalizability to modern ransomware, DDoS variants, and encrypted C2 traffic has not been empirically established and is discussed in Section 10. Behavioral indicators associated with ransomware propagation and lateral-movement patterns are detected in the Neris scenario; direct evaluation against post-2020 ransomware families remains Future Work §11.
All experiments were conducted in a reproducible Vagrant/VirtualBox environment running on a consumer laptop, ensuring result accessibility and independent reproducibility.
This work was developed through a structured multi-model AI collaboration methodology — the Consejo de Sabios (Council of Wise Men), a multi-LLM peer review methodology — documented in Section 6. Test-Driven Hardening (TDH) is proposed as a methodology for building security-critical distributed systems, emerging directly from this collaboration. ML Defender is released under the MIT license.
1 Introduction
The threat is not abstract. On March 5, 2023, the Hospital Clínic de Barcelona suffered a ransomware attack that paralyzed its emergency services, laboratory, and pharmacy, encrypted 4.5 TB of patient data, and prompted a ransom demand of 4.5 million USD — which the institution refused to pay [8]. That same day, the author of this work was leaving a regional hospital in Extremadura after receiving treatment for porphyria — a rare metabolic disease that his local hospital had diagnosed and managed. Watching the news, the thought was immediate and personal: if it can happen to a major teaching hospital, it can happen to mine.
This was not the first time that the author had witnessed the human cost of ransomware firsthand. Years earlier, the collapse of a close friend’s small business had been caused by an encryption attack that locked every document and file in his office. The author lacked the technical skills to help. That failure left a debt.
These two events — the friend’s destroyed business, the hospital on television — converged into a single design constraint: the solution had to be open-source, it had to run on commodity hardware, and it had to act in real time. Not forensics. Not post-mortem analysis. Interception.
The analogy that emerged from this motivation is biological. Network traffic behaves like a circulatory system: datagrams are the bloodstream, and malicious flows are pathogens. A useful defense must classify traffic continuously and in real time — capturing the source IP of the attacker and blocking it at the firewall before damage propagates. This framing shaped every architectural decision in ML Defender.
This work therefore addresses the following research question: can an embedded machine-learning network detection and response system running entirely on commodity hardware provide meaningful real-time protection for organizations lacking enterprise security infrastructure?
A note on terminology.
ML Defender (aRGus NDR) implements Network Detection and Response (NDR) capabilities: real-time capture, ML-based classification, and automated blocking at the network layer. Future work toward full Endpoint Detection and Response (EDR) functionality via a planned lightweight endpoint agent (§11, FEAT-EDR-1) will be reflected in a corresponding rename when that capability is delivered.
The gap.
Existing open-source NIDS — most notably Snort and Suricata [11] — operate primarily on signature matching and rule-based heuristics. While effective against known attack patterns, they offer limited resilience against novel ransomware variants and volumetric DDoS campaigns. Machine learning-based NIDS have been extensively studied [4, 12], but the gap between research prototype and deployable system remains wide: most published systems are evaluated on benchmark datasets without accompanying implementation, require GPU infrastructure, or assume dedicated security operations teams. Healthcare retained its position as the most targeted sector in 2025, accounting for 22% of disclosed ransomware attacks, with an average breach cost of 7.42M USD [3, 7].
Our approach.
ML Defender (aRGus NDR) addresses this gap with three design principles: (1) embedded inference — all ML classification runs in-process at sub-microsecond latency; (2) commodity deployment — the full six-component pipeline runs on bare-metal Linux hardware costing approximately 150–200 USD; (3) active response — upon detection, the attacker’s IP is blocked immediately at the firewall via ipset/iptables.
Beyond the technical system, this paper documents a second contribution: the Consejo de Sabios methodology for human-AI collaborative engineering.
2 Background and Related Work
Signature-based NIDS.
Snort [13] and Suricata [11] apply rule-based pattern matching against packet payloads and flow metadata. Their fundamental limitation is reactive: detection requires a prior known signature, a window that historically ranges from hours to weeks.
ML-based NIDS.
Anderson & McGrew [1] demonstrated the feasibility of contextual flow data for encrypted malware classification. Mirsky et al. [10] proposed Kitsune, an ensemble of autoencoders for unsupervised anomaly detection operating on flow statistics. The CTU-13 dataset [6], which ML Defender uses for evaluation, has become a standard benchmark for botnet traffic detection. Despite this body of work, a persistent gap exists between academic prototype and production system.
Flow-based ML NIDS.
ML Defender operates exclusively at the flow level, extracting 28 behavioral features per flow without examining payload content — ensuring that payload encryption does not compromise detection capability.
eBPF and XDP-based network monitoring.
Recent work has explored eBPF/XDP as a foundation for high-performance network security. ML Defender follows the XDP-capture-plus-userspace-inference architecture: XDP provides zero-copy packet capture before the kernel networking stack, while all ML inference executes in userspace at sub-microsecond latency. To our knowledge, open-source projects combining eBPF capture with userspace ML inference remain rare; most production-grade alternatives either offload inference to the cloud or require GPU hardware.
Embedded inference.
Commercial EDR vendors (CrowdStrike, SentinelOne, Carbon Black) implement proprietary embedded ML pipelines, but neither their architectures nor their training data are publicly available. ML Defender demonstrates that embedded Random Forest inference achieves sub-microsecond latency without requiring GPU acceleration or cloud offload.
Human-AI collaborative engineering.
Structured methodologies for multi-model peer review have not, to our knowledge, been formally documented in the systems security literature. Section 6 describes the Consejo de Sabios methodology.
AI-native security reasoning and the evolving threat landscape.
This paper was written and submitted in April 2026, concurrent with the announcement of Anthropic’s Project Glasswing [2], which demonstrated that AI models can autonomously identify and chain kernel-level vulnerabilities — including local privilege escalation to root in Linux — at a scale and speed previously requiring specialized human expertise. These results represent a shift in the threat landscape: AI-augmented offensive capabilities are no longer theoretical. This directly motivates the explicit kernel security boundary axiom in §3.3: aRGus NDR assumes the kernel as a potentially compromised boundary and shifts its trust anchor to verifiable network behavioral patterns. The network remains an observable chokepoint even when the host is not. The hardened deployment variants ADR-030 (AppArmor) and ADR-031 (seL4) documented in §3 are a direct architectural response to this trajectory.
3 Threat Model
3.1 Adversary Capabilities
The system assumes an adversary capable of launching attacks from external hosts or compromised internal machines, including: botnet C2 communication, DDoS (including amplification attacks), ransomware propagation via SMB lateral movement, and network reconnaissance.
3.2 Adversary Limitations
The system assumes the attacker cannot: fully mimic benign traffic distributions across all monitored features simultaneously; compromise the ML Defender host; manipulate feature extraction; or tamper with HMAC-SHA256 protected event logs.
3.3 Kernel Security Boundary (Explicit Axiom)
ML Defender assumes the host kernel is not actively compromised at deployment time. This is a declared scope boundary, not an oversight.
An adversary with root access or kernel-level persistence — via a rootkit, a compromised eBPF JIT, or memory-resident malware — can bypass userspace detection mechanisms regardless of their sophistication. This is a fundamental property of all userspace security software, not a specific weakness of ML Defender.
This assumption is consistent with the threat model of the target organizations (hospitals, schools, municipalities), where the realistic adversary is an opportunistic network attacker exploiting known vulnerabilities, not a state-level actor with pre-established kernel-level persistence. Against this realistic threat model, ML Defender’s network detection layer remains effective: lateral movement, C2 communication, and data exfiltration must traverse the network regardless of kernel state. The network is the choke point; it remains observable even when the host is not.
Hardened deployment variants addressing kernel security are documented as future work: ADR-030 (AppArmor-enforcing) and ADR-031 (seL4 microkernel), both described in §3.
3.4 Detection Surface
Detection occurs at the network flow level. ML Defender focuses on connection frequency, port diversity, TCP flag distributions, packet inter-arrival timing, flow burst behavior, and destination diversity — features detectable even when payloads are encrypted.
3.5 Out-of-Scope Attacks
Application-layer attacks generating statistically normal traffic; low-and-slow attacks within benign statistical distributions; insider threats; attacks targeting the ML Defender host itself. Future work in Section 11 addresses some of these through expanded detection vectors (FEAT-NET-1, FEAT-AUTH-1).
Physical and removable-media vectors (conscious design decision).
ML Defender does not monitor file system activity, removable storage, or USB-borne payloads. This is an intentional architectural boundary, not an oversight.
The guiding principle of the pipeline is network surveillance: every component — eBPF/XDP capture, flow-level feature extraction, ML-based classification, firewall response — operates on network traffic. File integrity and endpoint state are out of the monitored surface by design.
Removable-media threats are addressed at the organizational and hardware layer: USB ports in the DMZ should be physically or firmware-disabled by the IT team before deployment; internal policy should prohibit removable media on monitored hosts. This is standard practice under frameworks such as CIS Controls v8 [5] and is the responsibility of the deploying organization, not of an NDR system.
For organizations that require file integrity monitoring on top of network detection, ML Defender is designed to operate in complementary mode alongside battle-tested open-source tools such as Wazuh [15] and NIST-aligned scanners. Wazuh provides HIDS capabilities (file integrity monitoring, log analysis, rootkit detection) that cover precisely the endpoint surface that ML Defender consciously excludes. The two systems are architecturally orthogonal: ML Defender defends the network perimeter; Wazuh defends the host state. Together they constitute a layered defense appropriate for the target organizations. Integration via raw TCP event streaming is described as Future Work (§4.6).
4 Architecture
4.1 Pipeline Overview
Six components connected via ZeroMQ and Protocol Buffers:
  1. 1.
    eBPF/XDP Sniffer — zero-copy packet capture at the XDP hook before kernel networking stack.
  2. 2.
    Ring Consumer — reconstructs bidirectional flow state via ShardedFlowManager, computes 28 of 40 planned features. Eleven features use MISSING_FEATURE_SENTINEL = -9999.0f.
  3. 3.
    Fast Detector — rule-based heuristic engine. FPR on purely benign traffic: 6.61% (§8.6).
  4. 4.
[…]
  5. 5.
    RAG Subsystem (rag-ingester + rag-security) — rag-ingester ingests detection events into FAISS + SQLite. rag-security exposes a natural language query interface via local TinyLlama — deliberately local to avoid exfiltrating live network metadata to cloud-hosted LLMs.
  6. 6.
    Firewall ACL Agent — applies ipset/iptables rules on confirmed detection events.
Figure 1 illustrates the end-to-end flow from packet capture to automated firewall response.
NetworkTrafficeBPF/XDPSnifferRing Consumer(Features)FastDetectorMLDetectorFirewallACL Agentipset/iptablesrag-ingesterrag-securityetcd-server(Config)ZMQZMQZMQZMQeventsFAISSJSON config $\max(s_{\mathrm{fast}},\;s_{\mathrm{ml}})$
Figure 1: ML Defender end-to-end pipeline. Six components communicate over ZeroMQ with Protocol Buffers serialization and ChaCha20-Poly1305 encryption. etcd-server distributes JSON configuration at startup. The dual-score Maximum Threat Wins policy (Eq. 1) selects the arithmetic maximum of Fast Detector and ML Detector scores; the RAG subsystem provides semantic observability over the live event stream.
4.2 Three Functions of the Engine
Function 1 — Real-time detection, classification, and response. End-to-end from packet capture to firewall block in milliseconds.
Function 2 — Reliable dataset generation. ML Detector and Firewall ACL Agent produce structured HMAC-verified CSV logs continuously, forming the basis for retraining future ensemble models.
Function 3 — Real-time observability (the digital microscope). The RAG subsystem provides a semantic query interface over the live event stream.
4.3 Dual-Score Detection
$score_{\text{final}}=\max\!\left(score_{\text{fast}},\;score_{\text{ml}}\right)$ (1)
This is the arithmetic maximum over two continuous scores in $[0,1]$, not a logical OR over binary decisions. Both detectors output continuous confidence scores; the maximum operator selects the higher-confidence assessment. The ML Detector’s primary role is false positive suppression: the Fast Detector generates $\text{FPR}=6.61\%$ on purely benign traffic; the ML Detector reduces real production blocks to zero on the same corpus (approximately 500-fold reduction). Alternative AND-based consensus policies are planned (ADR-007, §11).
4.4 Transport and Serialization
ZeroMQ for asynchronous message passing; Protocol Buffers (proto3) for serialization; etcd for configuration distribution; ChaCha20-Poly1305 for inter-component encryption (selected for performance on hardware without AES-NI, including ARMv8); HMAC-SHA256 for CSV log integrity.
4.5 Deployment Model
The reference deployment targets a single bare-metal Linux node (kernel $\geq$ 5.8) with two NICs: one connected to the monitored network segment, one to the management network. The full pipeline runs on commodity hardware costing approximately 150–200 USD. Validated in Vagrant/VirtualBox; bare-metal characterization is Future Work (§11).
Figure 2 illustrates a representative high-availability deployment in a hospital environment, with a primary and a warm-standby node protecting two internal network segments.
Internet /External ThreatsML DefenderPrimary Node\$150–200 HWML DefenderWarm Standby\$150–200 HWheartbeatClinicalSystemsAdminNetworkMedical IoTDevicesManagementNetworketcd-server+ RAG queryNIC 1 (monitored)NIC 2 (mgmt) failover promotes ipset block on detection
Figure 2: Representative high-availability deployment of ML Defender in a hospital environment. A primary node monitors all inbound and inter-segment traffic via NIC 1 (monitored); a warm-standby node receives heartbeat signals and promotes automatically on primary failure. Both nodes run the full six-component pipeline on commodity hardware. etcd-server and the RAG query interface are co-located on the primary node. Firewall ACL Agent topology: scored Protocol Buffers events flow from ml-detector to one or more firewall-acl-agent instances, which apply ipset/iptables rules at the relevant boundary. A single gateway node requires one agent; each independent segment with its own firewall requires its own agent instance. etcd-server is shown in single-node mode (current implementation); a distributed HA etcd cluster is in the project backlog. Peer-to-peer seed negotiation is under design.
4.6 Integration Philosophy: Composability over Monolithism
ML Defender is designed to operate as a composable component within a broader security ecosystem, not as a monolithic solution claiming to cover every threat vector. The pipeline’s guiding principle — network surveillance — is complemented, not replaced, by tools that cover orthogonal surfaces.
Protocol design constraint.
All external integrations must use the same transport stack as the pipeline itself: raw TCP sockets for transport, Protocol Buffers for serialization, and ChaCha20-Poly1305 for encryption. HTTP, Kafka, WebSocket, and broker-based architectures are explicitly out of scope. This constraint is not arbitrary. Four independent arguments support it:
  1. 1.
    Deterministic latency. HTTP and Kafka introduce non-deterministic scheduling jitter incompatible with sub-10 ms response requirements. Raw TCP over ZeroMQ provides predictable, bounded delivery latency.
  2. 2.
    Reduced attack surface. HTTP parsers have historically been a primary source of remotely exploitable CVEs (buffer overflows, request smuggling, header injection). Raw TCP with Protocol Buffers framing eliminates this parser surface by more than 90%.
  3. 3.
    No broker, no single point of failure. Kafka and Redis require dedicated broker infrastructure incompatible with single-host deployments at the 150–200 USD target. Every additional process is a failure domain; the pipeline’s ZeroMQ topology has no mandatory broker.
  4. 4.
    Minimal footprint. Eliminating HTTP, Kafka, and WebSocket transport removes librdkafka, libcurl, and boost.asio from the dependency tree — a non-trivial reduction in binary size and attack surface for resource-constrained deployment targets.
FEAT-INT-1: Wazuh and NIST scanner integration (planned).
A planned integration layer will allow Wazuh agents [15] and NIST-aligned file integrity scanners deployed on DMZ hosts to emit events directly into the ML Defender pipeline via raw TCP sockets. Events are serialized as Protocol Buffers messages at the source and ingested by rag-ingester via the existing ZeroMQ transport. The Noise Protocol IK handshake (ADR-024) provides mutual authentication between the Wazuh agent adapter and rag-ingester.
The motivation is graph quality: correlating network anomalies detected by ml-detector and firewall-acl-agent with file integrity violations reported by Wazuh on the same host produces composite attack patterns that neither system observes independently. A ransomware attack in its lateral-movement phase typically exhibits simultaneous network spread and file encryption activity; correlation in the RAG subsystem surfaces this composite signature. This integration requires ADR-028 (RAG Ingestion Trust Model) to be approved before implementation.
4.7 Service Coordination (etcd-server)
Fulfills component registration, JSON config distribution, and ChaCha20 seed provision. ChaCha20 seed distribution via etcd is not recommended for production use — peer-to-peer seed negotiation is under design. etcd is currently deployed in single-node mode (see Figure 2); a distributed HA etcd cluster (3- or 5-node consensus) is in the project backlog and will be required for true high-availability deployments. The single-node constraint is the only architectural dependency preventing fully symmetric active/standby failover.
4.8 Fast Detector Heuristics (DEBT-FD-001)
Four heuristics in a 10-second sliding window:
  • External IP velocity: THRESHOLD_IP_VELOCITY = 5
  • SMB connection count: THRESHOLD_SMB_CONNS = 3
  • Port diversity: THRESHOLD_PORT_SCAN = 10
  • TCP RST ratio: THRESHOLD_RST_RATIO = 0.20
All thresholds are compile-time constexpr constants — DEBT-FD-001, scheduled for JSON migration in PHASE 2 (ADR-006).
5 Implementation
5.1 Language and Build System
C++20. CMake with four build profiles: production, debug, TSAN, ASAN.
5.2 eBPF/XDP Packet Capture
XDP hook attachment before SKB allocation. Zero-copy. All parameters loaded from sniffer/config/sniffer.json at startup.
5.3 Feature Extraction
ShardedFlowManager partitions flow tracking across shards, eliminating the ISSUE-003 thread-isolation bug (89 of 142 features silently lost in the original single-threaded FlowManager). 28 of 40 features computed; 11 use MISSING_FEATURE_SENTINEL = -9999.0f.
Sentinel value selection rationale.
Split-domain analysis confirmed all Random Forest thresholds lie within $[0.0,5.1]$. $9999.0\text{f}$ lies entirely outside this domain, routing deterministically to the left child of every split — non-informative, not influencing classification. quiet_NaN() was rejected (undefined comparison behavior across compilers); $0.5\text{f}$ was rejected (within domain, spurious classification influence). One feature uses $0.5\text{f}$ as a semantic sentinel for TCP half-open state (deliberate, within-domain meaningful default).
5.4 Embedded Random Forest and Synthetic Training Data
Three classifiers (Ransomware, DDoS, Traffic) compiled as C++20 data structures. Internal threats classifier via ONNX Runtime — deliberate to measure latency differential and footprint trade-off. Models trained with scikit-learn, transpiled to C++20 via custom script; zero Python or external runtime dependency.
Training/evaluation separation: all classifiers trained on synthetic data; CTU-13 Neris held out entirely for evaluation.
Synthetic Dataset Methodology.
Informed by CTU-13 (flow statistics), CIC-IDS2017 (protocol ratios), and MAWI (benign traffic). Four bias mitigation strategies: IP Space Partitioning (10.0.0.0/8 training vs 147.32.0.0/16 evaluation), Temporal Decoupling, Protocol Ratio Calibration, Feature Boundary Validation.
Known Limitations.
Reflects attack patterns circa 2011–2017. Temporal dynamics of real C2 channels simplified to fixed-window aggregates. Zero-day variants may evade classification.
5.5 Cryptographic Transport
ChaCha20-Poly1305 + LZ4 + HMAC-SHA256. Crypto-transport layer extracted as an independent library.
5.6 Plugin Architecture (ADR-023, ADR-024)
ML Defender implements a multi-layer plugin architecture (ADR-023) enabling runtime extension of pipeline components without modifying core detection logic. Plugins are loaded via dlopen/dlsym with a pure C ABI (PLUGIN_API_VERSION = 1), providing ABI stability across compiler versions and ensuring binary compatibility with future releases.
MessageContext and Graceful Degradation.
PHASE 2a (DAY 105) introduces MessageContext — a typed C struct passed to the optional plugin_process_message() symbol. Plugins that do not export this symbol are silently skipped (Graceful Degradation Policy D1), preserving backward compatibility with PHASE 1 plugins that implement only plugin_process_packet(). Post-invocation validation (D8) performs a byte-wise snapshot comparison of all read-only fields, detecting any plugin attempt to modify ownership-restricted data. The trust boundary is explicit: plugins are treated as untrusted code (D7); the final blocking decision remains exclusively in the core (ADR-012).
Threat Model Integration (ADR-023 D9).
The plugin system’s Trusted Computing Base (TCB) is limited to PluginLoader and CryptoTransport. Plugin code has no access to cryptographic keys, seed material, or internal transport state. The MLD_DEV_MODE escape hatch is restricted to Debug builds compiled with MLD_ALLOW_DEV_MODE (D10), ensuring fail-closed behavior in all production builds.
Gate TEST-INTEG-4a.
Integration of the plugin loader into firewall-acl-agent (PLUGIN-LOADER-FW) was validated by gate TEST-INTEG-4a: plugin_process_message() invoked on a live MessageContext, post-invocation invariants verified, result_code = 0 confirmed, and CryptoTransport decryption path verified unmodified by diff.
Forward Compatibility: Dynamic Group Key Agreement (ADR-024).
The MessageContext struct reserves 60 bytes (reserved[60]) for forward compatibility with ADR-024, which specifies Dynamic Group Key Agreement using the Noise Protocol Framework (Noise_IKpsk3 or Noise_KK pattern). ADR-024 design is approved; open questions OQ-5 through OQ-8 (static key revocation, session continuity during rotation, replay protection, and ARMv8 performance characterization) are under resolution. Implementation is scheduled post-arXiv submission.
5.7 Trace Correlation
trace_id = SHA-256(src_ip, dst_ip, canonical_attack_type, timestamp_bucket). TraceIdPolicy: ransomware=60s, DDoS=10s, SSH brute=30s. Validated: 46/46 unit tests passing.
5.8 Testing and Validation
70 tests total: crypto (3), etcd-hmac (12), ml-detector (9), trace_id (46). 100% pass rate. Integration gate TEST-INTEG-4a (plugin MessageContext invocation, D8 post-invocation validation) confirmed on DAY 105. F1 ground truth persisted to docs/experiments/f1_replay_log.csv.
5.9 Development Environment
Vagrant/VirtualBox. Host: MacBook Pro Intel Core i9 2.5 GHz, 32 GB DDR4. Guests: defender (Ubuntu 24.04 LTS, Linux 6.x, dual vNIC) + client (traffic generator).
6 The Consejo de Sabios: A Methodology for Human-AI Collaborative Engineering
6.1 Motivation
The development of ML Defender spans kernel programming (eBPF/XDP), distributed systems (ZeroMQ, etcd), cryptography, machine learning, systems security, and academic writing. The author operated as an independent researcher without institutional affiliation, without a team, and without access to a formal peer review process. The solution was to construct one.
6.2 The Methodology
Seven large language models served as co-reviewers across all development phases: Claude (Anthropic), Grok (xAI), ChatGPT (OpenAI), DeepSeek, Qwen (Alibaba), Gemini (Google), Parallel.ai.
These models were engaged as intellectual peers — presented with architectural proposals, asked to identify failure modes, invited to challenge assumptions. The methodology operates on a principle of structured disagreement: the Council does not average outputs. It introduces adversarial validation as a mechanism to expose hidden assumptions and validate architectural decisions. Convergent recommendations increase confidence; divergent recommendations mandate deeper investigation before any decision is closed.
6.3 How It Worked in Practice
Convergent recommendations across models increased confidence. Divergent recommendations triggered deeper investigation. Decisions such as the $9999.0\text{f}$ sentinel value, the trace_id correlation architecture, and the ShardedFlowManager design emerged from or were refined through these dialogues.
6.4 Test Driven Hardening
Test-Driven Hardening (TDH) is proposed as a methodology for building security-critical distributed systems in which correctness properties are protocol-level, not component-level, and therefore undetectable by static analysis or isolated unit tests. A standalone reference repository documenting the methodology and representative case studies is maintained at https://github.com/alonsoir/test-driven-hardening.
TDH methodology: (1) Witness — characterize anomalous behavior. (2) Demonstration test — produce a minimal test that fails and exposes the mechanism. (3) Independent solution — each model proposes a fix against the shared contract. (4) Human arbitration — author selects, weighing correctness, maintainability, and coherence.
TDH differs from TDD: the test is written before the fix, as a proof of failure — not before the feature, as a specification.
6.5 HKDF Context Symmetry: A Pedagogical Case Study in Test-Driven Hardening
A subtle but critical defect discovered during the cryptographic integration sprint of ML Defender illustrates a class of correctness errors that are invisible to static analysis yet detectable through end-to-end protocol testing. We document it here as a concrete instance of the Test-Driven Hardening methodology described in §6.
The defect.
The CryptoTransport layer derives symmetric keys using HKDF-SHA256 [9] with a context string (info parameter) that encodes the identity of the communicating parties. The original implementation parameterised this context by component name:
Listing 1: Incorrect HKDF context — component-scoped
// WRONG: each component uses its own name as HKDF context.
// Sniffer derives K_sniffer; ml-detector derives K_mldetector.
// K_sniffer != K_mldetector -> MAC authentication failure on every channel.
const std::string ctx = "ml-defender:" + component_name
+ ":" + version;
Because each component used its own name as the HKDF context, transmitter and receiver independently derived distinct subkeys — causing MAC authentication failure on every cross-component channel. The defect was latent in unit tests, which exercised each component in isolation using loopback (same context on both sides), where MAC verification trivially succeeds. It only manifested when two components with different names attempted to communicate end-to-end.
To make the causal chain explicit: an incorrect context string causes the transmitter and receiver to derive distinct subkeys, resulting in MAC verification failure at the receiver. A correct, symmetric context string produces identical derived keys on both sides, enabling successful authentication. This class of error is undetectable without end-to-end protocol validation.
The correct parameterisation encodes directionality:
Listing 2: Correct HKDF context — channel-scoped
// RIGHT: tx and rx derive distinct subkeys
const std::string ctx_tx = "ml-defender:" + component_name
+ ":" + version + ":tx";
const std::string ctx_rx = "ml-defender:" + component_name
+ ":" + version + ":rx";
The canonical definitions were consolidated in crypto_transport/include/crypto_transport/contexts.hpp, providing a single source of truth for all six pipeline components. This design avoids semantic overloading of data structures across layers — a common source of subtle security and correctness bugs, as demonstrated by the defect described above.
Why static analysis cannot catch this.
Both the erroneous and correct context strings share the type std::string. The C++ type system offers no mechanism to distinguish a transmit context from a receive context at the type level without a dedicated wrapper type — an abstraction whose overhead is not warranted in a resource-constrained deployment target. The compiler, sanitisers (ASan, TSan), and clang-tidy all accepted both formulations without warning. This is a semantic correctness property, not a syntactic one.
Detection via TEST-INTEG-3.
The defect was exposed by TEST-INTEG-3, an integration test exercising the full sniffer $\to$ ml-detector cryptographic channel with an intentional regression: the test temporarily restores the component-scoped context on one side and asserts that decryption fails with a MAC authentication error. Under the correct implementation, this regression triggers std::terminate() via the fail-closed handler; under the defective implementation, decryption succeeds — and the test fails accordingly.
The key insight is that TEST-INTEG-3 does not test encryption or decryption in isolation. It tests the protocol invariant: that a message encrypted on the transmit path is decryptable only on the corresponding receive path, and that any violation of this invariant is immediately fatal. No unit test of CryptoTransport in isolation could have established this property.
Lessons for cryptographic software.
This case study yields three transferable observations:
  1. 1.
    Cryptographic correctness is a protocol property. Components that are individually correct may be collectively insecure. End-to-end tests exercising full message round-trips across trust boundaries are the minimal viable test surface for authenticated encryption schemes.
  2. 2.
    Intentional regression tests are first-class citizens. A test that asserts failure under a known-bad configuration is as valuable as one that asserts success under a known-good one. The TDH methodology mandates both polarities.
  3. 3.
    Single source of truth for cryptographic parameters. Distributing context strings across six main.cpp files created a silent coordination hazard. Centralising them in a dedicated header eliminated an entire class of latent divergence.
The defect was discovered on Day 97 of the development timeline, three days after CryptoTransport was integrated across all pipeline components. Its detection required no external audit: the TDH test suite caught it during routine regression testing on the feature/bare-metal-arxiv branch (now superseded by feature/plugin-crypto).
6.6 Attribution
The author remained the final arbiter of all decisions; all empirical validation was performed by the human author. The seven models are credited as intellectual co-contributors in the project’s commit history, ADRs, and this paper. In the context of arXiv submission, the models are acknowledged in §Acknowledgments; formal co-authorship conventions for AI systems remain an evolving area, and this paper adopts the most transparent available approach consistent with emerging ACM and IEEE norms on AI contribution disclosure.
6.7 Why This Matters
The Consejo de Sabios democratizes access to deep technical peer review for independent researchers without institutional affiliation — parallel to ML Defender’s mission of democratizing enterprise-grade security for organizations without enterprise budgets.
6.8 Limitations
Models cannot run code or observe system behavior directly. Models can hallucinate — several incorrect technical suggestions were identified and rejected. The methodology augments human judgment; it does not replace it.
7 Formal System Model
7.1 Network Traffic Representation
Packets $P=\{p_{1},\dots,p_{n}\}$ grouped into bidirectional flows by canonical five-tuple $(src,dst,sport,dport,proto)$ via ShardedFlowManager $\Phi:P\rightarrow F$. Flows are bidirectional five-tuple aggregates; behavioral features are computed over a 10-second sliding window as configured in sniffer.json. Flows are closed after an inactivity timeout as defined by the sliding window expiry.
7.2 Feature Extraction
$x_{f}=g(P_{f},W),\quad x_{f}\in\mathbb{R}^{28}$ (2)
7.3 Fast Detector
$H(x_{f})=\max_{j}h_{j}(x_{f}),\quad h_{j}(x_{f})\in\{0,1\}$ (3)
7.4 Machine Learning Detector
$M(x_{f})=\frac{1}{T}\sum_{k=1}^{T}T_{k}(x_{f})$ (4)
7.5 Detection Policy
$S(x_{f})=\max\!\left(H(x_{f}),\;M(x_{f})\right)\geq\theta$ (5)
Thresholds: $\theta_{\text{ransomware}}=0.85$, $\theta_{\text{DDoS}}=0.90$, $\theta_{\text{traffic}}=0.80$, $\theta_{\text{internal}}=0.85$.
7.6 Response Function
$R(f)=\begin{cases}\texttt{block}(src_{f})&\text{if }S(x_{f})\geq\theta\\ \texttt{allow}&\text{otherwise}\end{cases}$ (6)
Blocking is applied per source IP. Blocks are currently persistent for the session duration; TTL-bounded blocks with configurable expiry are Future Work (§11).
7.7 System Determinism
Given identical packet input $P$, the pipeline produces identical outputs. Determinism is a prerequisite for the reproducibility guarantees described in Section 13.
8 Evaluation
8.1 Experimental Setup
Table 1: Experimental hardware and software specifications.
Parameter Value
Host machine Apple MacBook Pro (Intel Core i9, 2.5 GHz)
Host RAM 32 GB DDR4
Guest OS Ubuntu 24.04 LTS (Linux kernel 6.x)
Virtualization VirtualBox 7.2, bridged NIC mode
VM vCPUs 4 (estimated $\sim$}2.4 GHz under VirtualBox)
VM RAM 4 GB (8 GB allocated for stress tests)
Network adapter Paravirtualized (VirtIO)
Virt. overhead Estimated 5–15% vs bare-metal
Pipeline config. Full 6-component, dual-VM
Methodological note on metrics.
calculate_f1_neris.py measures Fast Detector alerts ([FAST ALERT] in sniffer.log) against CTU-13 Neris ground truth IPs. The ML Detector’s attacks=N statistic reports flows exceeding production thresholds and triggering firewall blocks. Both are reported; F1=0.9985 is the primary system detection metric.
8.2 Dataset
CTU-13 Neris.
19,135 total flows; 646 malicious flows (host 147.32.84.165) constitute the TP ground truth. Captured 2011. Garcia et al. [6] document that the Neris botnet exhibits C2 communication patterns concentrated in a behavioral subset — specifically, flows matching IRC-based C2 channel signatures (connection bursts, SMB lateral movement, DNS query anomalies). The remaining flows represent background traffic on the infected host that do not exhibit malicious behavioral signatures and are therefore not counted as ground-truth positives. This distinction aligns with standard practice in NIDS evaluation [14, 10].
BigFlows.
A CTU-13 derived dataset containing 40,467 flows and 791,615 packets, used for high-throughput replay experiments. 38,064 flows were processed by ml-detector. Treated conservatively as probable benign — no official ground truth exists for this CTU-13 scenario.
Synthetic training data.
All classifiers trained exclusively on synthetic data; CTU-13 Neris held out entirely for evaluation.
8.3 Detection Performance
Table 2: Detection performance — Fast Detector pipeline on CTU-13 Neris.
Metric Value
F1 Score 0.9985
Precision 0.9969
Recall 1.0000
True Positives 646
False Positives 2
False Negatives 0
True Negatives 12,075
Total events processed 12,723
FPR (Neris evaluation) 0.0002%
ML Detector high-conf. attacks 12
FPR Fast Detector (bigFlows) 6.61%
FP reduction (Fast $\to$ ML) $\approx$}500-fold
F1=0.9985 stable across replay runs at total_events = 12,563 / 12,605 / 12,723 / 13,930.
8.4 False Positive Identification and Confusion Matrix
Table 3: False positive identification.
FP Source Destination Classification
FP-1 192.168.56.1 224.0.0.251 mDNS multicast (VirtualBox host-only)
FP-2 192.168.56.1 192.168.56.255 Broadcast (VirtualBox host-only)
Both are VirtualBox host-only adapter artifacts. Their absence in bare-metal deployments is a reasonable expectation, but has not been empirically verified (§11).
Confusion matrix.
$\text{Precision}=\frac{646}{648}=0.9969,\quad\text{Recall}=\frac{646}{646}=1.0000,\quad\text{F1}=0.9985$
8.5 Inference Latency
Table 4: Per-class inference latency.
Classifier Implementation Latency
[…]
Latency values are mean inference times measured via std::chrono::high_resolution_clock around the classify() call within the ml-detector process, excluding inter-component ZeroMQ transport latency, over $N\geq 1{,}000$ inference calls per classifier during the CTU-13 Neris replay.
8.6 Ablation Study
Config A — Fast Detector Only: 2,517 FP on 38,064 bigFlows (FPR=6.61%), Recall=1.0000 on Neris.
Config B — ML Detector Only (inferred): 5 detections on 38,064 bigFlows, all below production thresholds (max confidence 68.97%) — 0 real blocks. Neris: attacks=12.
Config C — Dual (deployed): F1=0.9985, FPR=0.0002%. Only Config C fully validated end-to-end.
8.7 Interpretation and Comparison with State of the Art
F1=0.9985 and Recall=1.0000 reflect the statistical separability of the Neris botnet scenario, not a claim about arbitrary threats. The synthetic training successfully captured the Neris behavioral signature; generalization to other families is an open empirical question.
Table 5: Indicative comparison with representative NIDS approaches. Caution: cross-paper comparison is indicative only — systems evaluated on different datasets, at different operating points, and under different experimental conditions. Direct experimental comparison was not performed.
System Dataset F1 Latency HW Block Notes
[…]
Snort N/A N/A 1–5 ms x86 Yes Rule-based, no ML
8.8 Ransomware Detection: Scope and Validation
What is validated.
Behavioral indicators associated with ransomware propagation and lateral-movement patterns are detected in the 2011 Neris botnet scenario: SMB lateral movement (ports 445/139 burst), high-rate connection attempts, DNS query anomalies.
What is not validated.
Direct evaluation against post-2020 ransomware families (LockBit, BlackCat/ALPHV, Cl0p) has not been performed. Modern ransomware uses encrypted C2, adaptive beaconing with jitter, staged delivery — potentially different statistical signatures.
Future work.
Direct evaluation against modern ransomware captures is the highest-priority evaluation expansion (§11), with Sebastian Garcia (CTU Prague) as potential collaborator.
8.9 Throughput Stress Test and Resource Utilization
Table 6: Stress test replay results — progressive Mbps escalation (DAY 87).
Run Dataset Req. Actual PPS Packets Failed Errors
1 Neris 10 Mbps 9.60 Mbps $\sim$}3,700 320,524 2,630 0
2 Neris 25 Mbps 11.28 Mbps $\sim$}4,300 320,524 2,630 0
3 bigFlows 50 Mbps 34.68 Mbps 9–10K 791,615 0 0
4 bigFlows 100 Mbps ($\times$3) 33.16 Mbps 9–11K 2,374,845 0 0
$\dagger$ Pre-existing pcap artifact — identical count at 10 and 25 Mbps, confirming origin in pcap structure, not pipeline saturation.
Table 7: Resource utilization during 100 Mbps bigFlows replay (DAY 87).
Component CPU% RES MEM% Notes
ml-detector $\sim$}315–320% 164 MB 2.0% 3.2 cores; multi-threaded
sniffer $\sim$}88–108% 55–57 MB 0.7% XDP + ring consumer
firewall-acl $\sim$}10–16% 101 MB 1.2% Stable throughout
rag-ingester $\sim$}12–20% 36–39 MB 0.5% FAISS indexing
etcd-server $\sim$}0.2% 10 MB 0.1% Minimal
rag-security idle 1.2 GB VIRT 15.0% TinyLlama loaded
Total $\sim$}65–73% $\sim$}1,278 MB $\sim$}23–30% idle
$\ddagger$ rag-security’s 1.2 GB virtual footprint is the TinyLlama model mapped at startup.
The throughput ceiling is VirtualBox NIC emulation ($\sim$}33–38 Mbps), not pipeline logic. Zero packet drops, zero deserialization errors across all 2,374,845 packets. Total pipeline RAM stable at $\sim$}1.28 GB (${\pm}18$ MB drift) throughout the 8-minute test.
9 Performance Model and Throughput Analysis
9.1 Pipeline Processing Model
Six stages operating concurrently via ZeroMQ; effective throughput bounded by the slowest stage:
$T_{\max}=\frac{1}{\max_{i}(L_{i})}$ (7)
9.2 Inference Cost
$\mathcal{O}(T\cdot D)$ — simple threshold comparisons on compact C++20 data structures fitting in CPU cache.
9.3 Throughput: Theoretical Bound and Measured Ceiling
Theoretical upper bound (single-core, ML inference only):
$T_{\text{theory}}\approx\frac{1}{0.5\;\mu\text{s}}=2\times 10^{6}\;\text{flows/sec}$ (8)
Measured virtualized ceiling (DAY 87).
The VirtIO NIC limits delivery to $\sim$}33–38 Mbps ($\sim$}9,000–11,000 pps). The observed throughput is limited by the virtualized NIC, not the pipeline. At this ceiling: ml-detector sustains $\sim$}3.2 cores; total pipeline RAM stable at $\sim$}1.28 GB; system CPU $\sim$}65–73% user, $\sim$}23–30% idle.
9.4 Queue Stability
Stability requires $\lambda<\mu$. Post-replay drain behavior provides direct empirical evidence: sniffer CPU drops to near-zero while ml-detector continues draining at full throughput for several minutes — confirming the ZeroMQ buffer absorbed the burst cleanly. errors=(deser:0, feat:0, inf:0) confirmed at all observation points.
9.5 Scalability
Vertical: additional CPU cores $\to$ additional parallel flow-processing threads, linear throughput improvement up to capture bottleneck.
Horizontal: multiple ML Defender nodes across network segments, each processing local traffic independently while contributing to distributed dataset generation.
10 Limitations
10.1 Evaluation Scope. Single botnet scenario (CTU-13 Neris, 2011). Modern ransomware communication patterns may differ substantially. Generalizability not empirically established.
10.2 Incomplete Feature Set. 28 of 40 features implemented; 11 use 9999.0-9999.0f sentinel. Impact of missing 12 features on broader traffic distributions unknown.
10.3 Synthetic Training Data. Reflects patterns circa 2011–2017. Zero-day variants may evade classification.
[…]
10.5 Cryptographic Seed Distribution. ChaCha20 seed via etcd not recommended for production. Peer-to-peer negotiation under design.
10.6 Single-Node Deployment. Distributed deployment not validated experimentally.
10.7 etcd High Availability. Single-node; failure disrupts registration and config distribution. HA planned.
10.8 RAG Index Capacity. FAISS event store finite capacity not mathematically characterized.
10.9 Virtualization Overhead. All latency and throughput measurements obtained under VirtualBox. Reported latencies are conservative upper bounds; reported throughput ($\sim$}33–38 Mbps) is a conservative lower bound imposed by NIC emulation. Bare-metal characterization remains future work (§11).
10.10 BigFlows Ground Truth. Treated as probable benign, not confirmed benign. FPR=0.0002% is a lower bound.
10.11 Threats to Validity.
  1. 1.
    Seed compromise. If seed.bin is exfiltrated, an adversary can derive all component subkeys and forge authenticated messages. Mitigations: strict file permissions (0600), in-memory zeroing via explicit_bzero, and planned peer-to-peer key agreement (ADR-024).
  2. 2.
    Single-host key assumption. All six components currently share a common seed family without distributed key agreement. Isolation between deployments relies on distinct seed provisioning per instance (ADR-021); formal distributed key agreement is post-arXiv work (ADR-024).
  3. 3.
    No hardware root of trust. The current design does not leverage TPM or secure enclave primitives. Physical access to the host enables seed extraction. Hardware-anchored key storage is identified as a future hardening target.
10.12 Kernel Security Boundary. ML Defender operates in userspace; eBPF/XDP programs execute in the kernel but are loaded by userspace. An adversary with pre-existing kernel-level access (rootkit, compromised eBPF JIT) can bypass the detection pipeline. This limitation is shared by all userspace NDR systems and is consistent with the declared scope boundary (§3.3). The realistic threat model for target organizations does not assume such adversaries: opportunistic ransomware operators attack via the network, not via pre-established kernel persistence. Network detection remains a valid and necessary defensive layer even under this constraint. Hardened variants (ADR-030, ADR-031) address this in future work (§3).
11 Future Work
11.1 Evaluation Corpus Expansion (P0). CTU-13 additional scenarios, modern ransomware captures, volumetric DDoS datasets. Contact initiated with Sebastian Garcia (CTU Prague) as potential arXiv endorser and collaborator.
11.2 Feature Completion and Importance Analysis. Complete remaining 12 real features. Feature importance analysis (Gini impurity reduction per feature from embedded RF).
11.3 Fast Detector JSON Migration (DEBT-FD-001). Scheduled for PHASE 2.
11.4 Consensus Detection Policy (ADR-007). AND-based policy requiring simultaneous agreement before autonomous blocking — JSON-configurable for operator selection.
11.5 Response Strategy Expansion. TTL-bounded blocks, traffic quarantine, honeypot redirection.
11.6 Cryptographic Hardening. Peer-to-peer ChaCha20 seed negotiation; etcd HA.
11.7 Runtime Reconfiguration (Hot-Reload). JSON parameter updates without restart.
11.8 Graph Attack Engine (ENT-2). Live attack graphs from CSV event data; STIX 2.1 export.
11.9 Fleet-Wide Observability. Fine-tuned LLM over anonymized aggregated data from all deployed nodes.
11.10 Test Driven Hardening as Independent Research. TDH generalizability, scaling with model capability, and measurable outcomes vs single-model consultation.
11.11 Bare-Metal Throughput Stress Testing (P1). Progressive tcpreplay at 100/250/500/1000 Mbps on physical NIC; CPU/memory/packet-drop monitoring per component. All throughput figures in this paper should be treated as conservative lower bounds.
11.12 Infiltration Vector Coverage. ML Defender currently detects ransomware propagation. Future work addresses infiltration: FEAT-NET-1 (DNS anomaly/DGA detection), FEAT-NET-2 (threat intelligence feeds), FEAT-AUTH-1 (authentication log ingestion), FEAT-AUTH-2 (behavioral auth anomaly), FEAT-EDR-1 (lightweight endpoint agent).
11.13 Enterprise Features. ENT-1 (Federated Threat Intelligence), ENT-2 (Attack Graph Generation), ENT-3 (P2P Seed Distribution), ENT-4 (Hot-Reload). MIT-licensed core remains free.
11.14 Ethical Deployment in Healthcare Environments. Not certified for FDA 21 CFR Part 820 or EU MDR 2017/745. Phased deployment: (1) 30-day monitoring mode with human approval; (2) TTL-bounded blocks with SOC notification; (3) full autonomous blocking after site-specific FPR validation.
11.15 Federated Distributed Intelligence (ENT-1 Extension). The current architecture processes all detection locally at each deployed node. A planned federated extension — not yet designed or implemented, and explicitly scoped to post-pipeline-completion milestones — would enable a coordinated fleet of ML Defender nodes to contribute anonymized telemetry to a central aggregation tier for continuous model retraining, with updated classifiers distributed back to the fleet as signed plugins (ADR-025). Figure 3 illustrates the intended architecture; implementation is contingent on full pipeline stabilization and will be reflected in a future revision of this paper.
Hospital AML DefenderSchool BML DefenderMunicipality CML DefenderHospital DML DefenderAggregation+ Training Tier(cloud / on-prem)Signed Plugins(ADR-025) anonymized telemetry retrained classifiers signed plugin distribution Planned — not yet implemented.All telemetry anonymized at source.Blocking decisions remain local.
Figure 3: Planned federated distributed intelligence architecture (not yet implemented). Each deployed ML Defender node operates fully autonomously; anonymized flow-level telemetry is contributed to a central aggregation tier for periodic retraining. Updated classifiers are packaged as Ed25519-signed plugins (ADR-025) and distributed to the fleet. All detection and response decisions remain local to each node; no live network metadata leaves the deployment site. This architecture is a post-pipeline-stabilization milestone; a future revision of this paper will document the implementation.
11.16 Wazuh/NIST Integration (FEAT-INT-1). Raw TCP streaming from Wazuh agents [15] and NIST-aligned scanners into rag-ingester for composite attack graph correlation. Protocol design, motivation, and architectural constraints documented in §4.6. Requires ADR-028 (RAG Ingestion Trust Model). Deferred to post-PHASE-2 pipeline stabilization.
11.17 Hardened Deployment Variants (ADR-030, ADR-031). Two hardened deployment variants are documented as architectural decisions, both targeting a stronger kernel security posture:
ADR-030 (aRGus-AppArmor-Hardened): Debian 13 + Linux 6.12 LTS + AppArmor enforcing mode. Full eBPF/XDP pipeline preserved. Compatible with ARM64 (Raspberry Pi 4/5) and x86-64. This is the planned production-hardened variant, providing mandatory access control against confused-deputy attacks on plugin loading (motivated by the analysis of Hugo Vázquez Caramés). Activation: post-PHASE-3 + hardware acquisition.
ADR-031 (aRGus-seL4-Genode): Pure research variant on the seL4 formally verified microkernel ($\sim$}12,000 lines of C verified in Isabelle/HOL). Linux runs as an unprivileged guest under Genode; XDP is infeasible in this model (no direct DMA access from guest), so libpcap is the mandatory capture fallback. The performance delta between eBPF/XDP and libpcap under seL4 is itself a measurable scientific contribution. Estimated overhead: 40–60%. Activation: post-ADR-030 + technical spike (2–3 weeks).
Both variants will be validated in Vagrant first, then on bare-metal. Results will be published with measured honesty; ADR-031 will not be merged into main regardless of outcome.
12 Conclusion
This paper presented ML Defender (aRGus NDR), an open-source network detection and response system built in C++20, designed from first principles to protect the organizations that need protection most and can afford it least.
[…]
Under progressive load testing, the pipeline processes up to $\sim$}34–38 Mbps — the ceiling imposed by VirtualBox NIC emulation — with zero packet drops, zero errors, and stable RAM ($\sim$}1.28 GB) across 2.37 million replayed packets. The bottleneck is the hypervisor, not the software.
These results cover one botnet scenario from 2011. They demonstrate architectural viability and effective detection within that scope. The limitations in Section 10 are the honest documentation of a system that does what it claims to do within the boundaries where it has been tested.
ML Defender was built because a friend lost his business to ransomware, and because a hospital in Barcelona was paralyzed by an attack on the same day its author was leaving a regional hospital in Extremadura — the hospital that had diagnosed his porphyria and saved his life. The design constraints that followed from those two events were not abstract: open-source, commodity hardware, real-time response. Not a forensic tool. Not a dashboard. A system that acts.
The Consejo de Sabios represents a new mode of engineering: one that gives a single independent researcher access to the kind of multi-expert peer review that has historically required institutional affiliation, research groups, and geographic proximity to centers of expertise. The Test Driven Hardening methodology that emerged from this collaboration has proven effective against some of the most subtle concurrency and memory bugs in the codebase.
Both contributions — the technical system and the collaborative methodology — are acts of democratization. The parallel is not accidental. It is the point.
For the hospital in Extremadura. For the friend whose business was broken by ransomware. For every small organization that has been told, implicitly or explicitly, that enterprise-grade security is not for them.
It is now.
13 Reproducibility and Artifact Availability
All experiments are reproducible from the public repository: https://github.com/alonsoir/argus
Dataset Setup.
# CTU-13 Neris (scenario 10) -- required for F1 validation
wget https://mcfp.felk.cvut.cz/publicDatasets/CTU-Malware-Capture-Botnet-42/\
botnet-capture-20110810-neris.pcap -O /vagrant/datasets/ctu13/neris.pcap
# Mirror: https://www.stratosphereips.org/datasets-ctu13
# CTU-13 bigFlows -- required for stress test
wget https://mcfp.felk.cvut.cz/publicDatasets/CTU-Malware-Capture-Botnet-42/\
bigFlows.pcap -O /vagrant/datasets/ctu13/bigFlows.pcap
F1 Evaluation.
make pipeline-stop && make logs-lab-clean && make pipeline-start && sleep 15
make test-replay-neris
python3 scripts/calculate_f1_neris.py \
/vagrant/logs/lab/sniffer.log --total-events 19135
Stress Test.
vagrant ssh client -c "sudo tcpreplay -i eth1 --mbps=100 --loop=3 \
/vagrant/datasets/ctu13/bigFlows.pcap"
Determinism.
Classifiers embedded in binary; no stochastic inference components; feature extraction deterministic. F1=0.9985 confirmed stable at total_events = 12,563 / 12,605 / 12,723 / 13,930.
Acknowledgments
This work was carried out independently, without institutional funding, without a research group, and without affiliation to any university or laboratory.
The author engaged seven large language models — Claude (Anthropic), Grok (xAI), ChatGPT (OpenAI), DeepSeek, Qwen (Alibaba), Gemini (Google), and Parallel.ai — as structured intellectual collaborators across 88 days of continuous development. Their contributions to architectural decisions, failure mode identification, debugging, test scenarios, and paper review are documented in the project’s commit history and ADRs. The Test Driven Hardening methodology documented in Section 6 emerged directly from this collaboration.
This paper discloses AI involvement openly and precisely. The author remained the final arbiter of all decisions; all empirical validation was performed by the human author.
The author wishes to acknowledge the friend whose business was destroyed by ransomware — who will recognize himself in these pages. And the Hospital Universitario de Badajoz, Extremadura, whose physicians diagnosed a rare disease and gave the author the years needed to build something worth building. Thank you Pedro Risco and Agustín Pijierro.
This work is dedicated to every hospital, school, and small organization that has ever been told that enterprise-grade security is beyond their reach.

### 19 Apr 2026 v3

@@
 Test-Driven Hardening (TDH) is proposed as a methodology for building security-critical distributed systems, emerging directly from this collaboration.
+[…]
 ML Defender is released under the MIT license.
@@
 Despite this body of work, a persistent gap exists between academic prototype and production system.
+Sommer and Paxson [16] identified fundamental reasons for this gap: the closed-world assumption of academic datasets, high base-rate asymmetry, and the absence of diversity in training distributions.
+ML Defender’s DAY 122 evaluation provides new empirical evidence for this thesis (§8.11).
 Flow-based ML NIDS.
@@
 Total pipeline RAM stable at $\sim$}1.28 GB (${\pm}18$ MB drift) throughout the 8-minute test.
+8.10 XGBoost Level-1 Classifier: In-Distribution Performance
+As part of PHASE 4 (ADR-026), a gradient-boosted classifier (XGBoost 3.2.0) was integrated as a hot-swappable plugin, replacing the Random Forest level-1 detector while preserving the Ed25519 signing chain and fail-closed std::terminate() contract of ADR-025.
+Experimental protocol.
+CIC-IDS-2017 [15] was partitioned by day to enforce temporal isolation: Tuesday, Thursday (morning and afternoon), and Friday (morning, DDoS, PortScan) constitute the training corpus; 20% of the training data was held out as a stratified validation set for threshold calibration and early stopping; Wednesday was sealed as a blind test set (MD5: bf0dd7e9...) and opened exactly once for final reporting.
+The threshold was calibrated exclusively on the validation set; no information from Wednesday influenced any modelling decision.
+Training details.
+23 flow-level features; scale_pos_weight=4.273=4.273 (1,303,148 benign / 304,974 attack); max_depth=6=6, eta=0.05=0.05, subsample=0.8=0.8; early stopping at round 724 of 1,000; val-AUCPR=0.99846=0.99846.
+Table 8: XGBoost level-1 performance on the validation set (in-distribution).
+Metric Value
+Precision (validation) 0.9945
+Recall (validation) 0.9818
+F1 (validation) 0.9881
+Calibrated threshold 0.8211
+[…]
+Best iteration 724 / 1,000
+Val-AUCPR 0.99846
+[…]
+8.11 XGBoost Level-1: Wednesday Out-of-Distribution Evaluation
+Result.
+[…]
+Table 9: Wednesday held-out threshold sweep (OOD). No operating point exists satisfying both medical gates simultaneously.
+Threshold Precision Recall FP/hour
+0.01 0.798 0.559 4,478
+0.10 0.776 0.202 1,841
+0.50 0.951 0.055 88
+0.82 (calibrated) 0.987 0.024 12
+0.90 0.985 0.017 8
+[…]
+Diagnosis: structural covariate shift by dataset design.
+Wednesday contains exclusively application-layer DoS attacks: DoS Hulk (n=231,073n=231{,}073; 91.5% of Wednesday attacks), DoS GoldenEye (n=10,293n=10{,}293), DoS Slowloris (n=5,796n=5{,}796), DoS Slowhttptest (n=5,499n=5{,}499), and Heartbleed ($n=11$).
+None of these attack types appear in any training CSV.
+CIC-IDS-2017 was designed with attack types segregated by day without cross-day repetition [15]; this design makes out-of-distribution generalization structurally impossible for any supervised classifier trained on the remaining days, regardless of algorithm or hyperparameter choices.
+DoS Hulk imitates high-volume HTTP traffic: its flow-level features overlap substantially with benign high-throughput flows.
+The model assigns a median probability of 0.0165 to Hulk flows — indistinguishable from benign traffic under any operationally viable threshold.
+Impossibility result.
+[…]
+Architectural implication.
+This result does not invalidate the in-distribution performance in §8.10.
+It establishes that static classifiers trained on academic benchmarks cannot be assumed to generalize to attack types absent from training.
+The aRGus architecture was designed for this reality: the XGBoost plugin is Ed25519-signed and hot-swappable (ADR-026).
+The loop enabling retraining from real adversarial traffic captured in the deployment environment is introduced as §3.
+The Wednesday evaluation report (wednesday_eval_report.json, MD5-sealed) is included in the repository as a permanent scientific artifact.
 9 Performance Model and Throughput Analysis
@@
 Hardened variants (ADR-030, ADR-031) address this in future work (§3).
+10.13 Structural Bias in Academic Datasets for Production NDR.
+The CIC-IDS-2017 temporal holdout evaluation (§8.11) demonstrates that academic benchmark datasets are structurally insufficient as the sole training source for production NDR classifiers.
+The day-specific attack segregation in CIC-IDS-2017 — and similar design choices in NSL-KDD, UNSW-NB15, and related corpora — creates a closed-world training distribution that does not reflect the open-world diversity of real network environments.
+This limitation is not specific to XGBoost, Random Forest, or any particular algorithm: it is a property of the data regime, as established by Sommer and Paxson [16] and quantified here with a complete threshold sweep.
+Synthetic data generated by language models (used for DDoS and ransomware classifiers in earlier phases) shares a related limitation: the generating model has not observed sufficient real adversarial traffic to reproduce the full statistical diversity of production attack flows.
+The only reliable path to production-grade classifiers for critical infrastructure NDR is training on traffic captured in the actual deployment environment, under controlled adversarial conditions.
+This motivates the Adversarial Capture-Retrain Loop (§3).
 11 Future Work
@@
 Activation: post-ADR-030 + technical spike (2–3 weeks).
-Both variants will be validated in Vagrant first, then on bare-metal. Results will be published with measured honesty; ADR-031 will not be merged into main regardless of outcome.
+Both variants will be validated in Vagrant first, then on bare-metal.
+Results will be published with measured honesty; ADR-031 will not be merged into main regardless of outcome.
+11.18 Adversarial Capture-Retrain Loop (ACRL).
+The out-of-distribution finding documented in §8.11 motivates a production retraining methodology that does not depend on academic benchmark datasets.
+The proposed Adversarial Capture-Retrain Loop (ACRL) operates as a closed cycle: (1) a red team agent (initially MITRE Caldera [11] for deterministic ATT&CK-mapped emulation; later a generative adversarial agent for adaptive evasion variants) executes attack techniques against an isolated network segment; (2) the aRGus pipeline captures all resulting traffic via eBPF/XDP, producing ground-truth-labelled flows; (3) the XGBoost plugin is retrained on the captured flows and re-evaluated against a held-out portion of the same session; (4) the retrained model is signed (Ed25519, ADR-025) and hot-swapped without pipeline restart (ADR-026).
+The loop addresses the closed-world limitation identified in §10 (10.13): each retraining cycle expands the training distribution toward the real attack surface of the deployment environment.
+The architectural prerequisite — hot-swappable signed plugins — is already implemented and validated.
+Minimum scientific validity requirements for the ACRL data source: (i) deterministic seed and versioned scenario for reproducibility; (ii) ground-truth labels at flow granularity, derived from red team execution logs; (iii) ATT&CK tactic coverage across at minimum three families (volumetric, protocol-layer, application-layer); (iv) network-valid traffic (correct TCP/IP headers, stateful handshakes); (v) isolated sandbox with no production traffic contamination.
+The term Adversarial Capture-Retrain Loop is introduced here as a named architectural pattern; related concepts include adversarial data flywheel and red team loop [16].
+Implementation is tracked as DEBT-PENTESTER-LOOP-001 in the project backlog.
 12 Conclusion
@@
 The limitations in Section 10 are the honest documentation of a system that does what it claims to do within the boundaries where it has been tested.
+A subsequent evaluation (DAY 122) integrated XGBoost 3.2.0 as a hot-swappable plugin and conducted a rigorous temporal holdout study on CIC-IDS-2017.
+The in-distribution results (Precision=0.9945=0.9945, Recall=0.9818=0.9818) confirm that the plugin architecture performs as designed.
+The out-of-distribution finding — that no threshold satisfies both medical gates on Wednesday data, because application-layer DoS attacks are structurally absent from the training days — is reported not as a failure but as a contribution: a quantified, empirically grounded argument for why production NDR classifiers must be trained on traffic from their actual deployment environment.
+The Adversarial Capture-Retrain Loop (§3) is the architectural response.
 ML Defender was built because a friend lost his business to ransomware, and because a hospital in Barcelona was paralyzed by an attack on the same day its author was leaving a regional hospital in Extremadura — the hospital that had diagnosed his porphyria and saved his life.